OpenStudy (anonymous):

Solve for x: e^2x=3x^2

OpenStudy (zzr0ck3r):

\[e^{2x}=3x^2\]?

OpenStudy (anonymous):

yes

OpenStudy (zzr0ck3r):

im not sure we can.

OpenStudy (anonymous):

using a graphing calculator?

OpenStudy (zzr0ck3r):

i could be wrong:) looks like mathstudent has something

OpenStudy (zzr0ck3r):

well you can get an approximate in a calculator solve(e^(2x)=3x^2,x) this is how you would do it on a ti-89

OpenStudy (anonymous):

I have a ti-84

OpenStudy (zzr0ck3r):

same thing should work

OpenStudy (zzr0ck3r):

lets see what @mathstudent55 says

OpenStudy (anonymous):

okay!

OpenStudy (mathstudent55):

\(e^{2x} = 3x^2 \) \(\ln (e^{2x}) = \ln(3x^2) \) \(2x = \ln3 + 2 \ln x \) \( 2x - 2 \ln x = \ln 3\) \( 2(x - \ln x) = \ln 3 \) \(x - \ln x = \dfrac{\ln 3}{2} \) I can't go any further.

OpenStudy (zzr0ck3r):

I think its transcendental and cant be solved. for an approximate http://www.wolframalpha.com/input/?i=approximate+e%5E%282x%29+%3D+3x%5E2

OpenStudy (zzr0ck3r):

http://en.wikipedia.org/wiki/Transcendental_equation read that @lily104

OpenStudy (anonymous):

thank you!

OpenStudy (amistre64):

\[e^{u}=\sum_0\frac1{n!}u^n\] \[e^{2x}=\sum_0\frac1{n!}(2x)^n\] \[e^{2x}=\sum_0\frac{2^n}{n!}x^n\] \[e^{2x}=1+\sum_1\frac{2^n}{n!}x^n\] \[e^{2x}=1+2x+\sum_2\frac{2^n}{n!}x^n\] \[\color{red}{3x^2} =1+2x+2x^2+\sum_3\frac{2^n}{n!}x^n\] \[0=1+2x+(2x^2\color{red}{-3x^2})+\sum_3\frac{2^n}{n!}x^n\] \[0=1+2x-x^2+\sum_3\frac{2^n}{n!}x^n\] \[x^2-2x+1=\sum_3\frac{2^n}{n!}x^n\]:)

OpenStudy (amistre64):

.. x^2-2x-1 that is